OpenStudy (anonymous):

-27/32 multiply 14/15 multiply by 3

hero (hero):

\[-\frac{27}{32} \times \frac{14}{15} \times 3\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cross multiply to get \[-9\times7\times3/16\times5=-189/80\]

hero (hero):

Okay, so... \[-\frac{27}{32} \times \frac{14}{15} \times \frac{3}{1}\] Multiply all the numerators (all the numbers at the top) Then multiply all the denominators (all the numbers at the bottom) Then simplify result if possible

OpenStudy (anonymous):

so 1134/480

hero (hero):

Actually you can change that into one big fraction, then simplify elements with common factors: \[-\frac{27 \times 14 \times 3}{32 \times 15 \times 1}\]

hero (hero):

For example, 27 and 15 is reducable by common factor 3 32 and 14 are reducable by common factor 2

OpenStudy (anonymous):

shouldndnt be 7/5

hero (hero):

If you reduce 27 and 15 by common factor 3, you divide both by 3 to get 9 and 5 If you reduce 32 and 14 by common factor 2, you divide both by 2 to get 16 and 7 Then you can re-write the fraction like so: \[\frac{9 \times 7 \times 3}{16 \times 5 \times 1}\]

OpenStudy (anonymous):

so 189/80

hero (hero):

forgot the negative, my bad.

OpenStudy (anonymous):

so -189/80

hero (hero):

Yes, correct

OpenStudy (anonymous):

thank you

hero (hero):

By the way, if you want to make a fraction like me, open equation editor and type frac{a}{b} to get: \[\frac{a}{b}\]

OpenStudy (anonymous):

ok dear..

OpenStudy (anonymous):

when i start a new question it does not give me that equation editor

hero (hero):

You can use the equation editor here to create it, then copy and paste the tex format (include the parentheses and brackets) into the question

OpenStudy (anonymous):

oh ok thanks